### v1

# Deep Context-Conditioned Anomaly Detection for Tabular Data

Abstract.
Anomaly detection is critical in domains such as cybersecurity and finance, especially when working with large-scale tabular data. Yet, unsupervised anomaly detection—where no labeled anomalies are available—remains a significant challenge. Although various deep learning methods have been proposed to model a dataset’s joint distribution, real-world tabular data often contain heterogeneous contexts (e.g., different users), making globally rare events normal under certain contexts. Consequently, relying on a single global distribution can overlook these contextual nuances, degrading detection performance. In this paper, we present a context-conditional anomaly detection framework tailored for tabular datasets. Our approach automatically identifies context features and models the conditional data distribution using a simple deep autoencoder. Extensive experiments on multiple tabular benchmark datasets demonstrate that our method outperforms state-of-the-art approaches, underscoring the importance of context in accurately distinguishing anomalous from normal instances.
Unsupervised Anomaly Detection, Deep Autoencoder
$c$s: Computing methodologies Anomaly detection
1. Introduction
Unsupervised anomaly detection (AD) is a critical field focused on identifying patterns that deviate from expected behavior, with applications spanning a wide range of tasks like security, fraud detection, and system reliability (Chandola et al., 2009; Aggarwal, 2017). The core challenge is not merely detecting new or unseen objects, but discerning whether such observations align with the learned notion of “normal,” or truly represent anomalies.
Recent advances in deep learning have driven a wave of neural network-based methods for unsupervised anomaly detection (Landauer et al., 2023; Pang et al., 2021b). Despite methodological differences, most deep approaches model the joint distribution of all features, implicitly assuming that data patterns are homogeneous across conditions. In practice, patterns often vary significantly with context. For example, in cybersecurity, a spike in data transfer may be benign for a user who routinely performs backups but suspicious for a database administrator. Ignoring such context can degrade detection accuracy.
To address this challenge, we argue probabilistically and demonstrate empirically the effectiveness of modeling conditional distributions over joint distributions for anomaly detection. We advocate for the context-conditional anomaly detection framework, which models conditional distributions to capture intra-context regularities while isolating inter-context variability, thereby enhancing detection performance without requiring separate models for each condition.
Selecting an appropriate context feature is a critical aspect of contextual learning. It is often not obvious what feature one should condition on when modeling the data. Indeed, the choice of conditioning variable can significantly influence model performance. To address this, we develop a bi-level optimization methodology to automatically identify the context features based on negative log-likelihood. Requiring only a small set of anomaly-free validation samples, our approach is simple, model-agnostic, and demonstrates consistent benefits across a wide range of datasets.
To empirically validate our claim, we leverage recent advancements in deep learning and advocate for the conditional Wasserstein autoencoder (CWAE), a conditional variant of the Wasserstein autoencoder (WAE) (Tolstikhin et al., 2019), to model conditional distributions for anomaly detection. We conduct experiments using several tabular AD benchmark datasets containing inherent contextual information (e.g., users, accounts, or hosts). We demonstrate that incorporating contextual information with the proposed CWAE model enables more robust anomaly detection across diverse and complex datasets, and offers significant advantages in improving detection accuracy.
Main contributions:
  • We promote a context-conditioned anomaly detection framework.
  • We propose the CWAE model to demonstrate the benefits of context-aware learning.
  • We introduce a principled method to select a suitable feature context feature for conditional modeling.
  • We demonstrate that conditional models can outperform both their non-contextual counterparts of the same architecture and state-of-the-art (SOTA) unconditional models.
2. Related Works
Anomaly detection has been a long-standing area of research, with early work dominated by statistical and distance-based approaches. Classical statistical methods such as Principal Component Analysis (PCA) (Shyu et al., 2003) and distance-based techniques like k-Nearest Neighbors (Liao and Vemuri, 2002) and clustering-based methods (Münz et al., 2007) aim to identify points that deviate from expected norms in the data distribution. While effective in lower dimensions, these techniques struggle in high-dimensional settings due to the “curse of dimensionality.”
More recent advances leverage deep neural networks to learn expressive representations for anomaly detection in complex data. Notable lines of work include one-class classification with deep feature embeddings (Ruff et al., 2018, 2021), autoencoder-based approaches that reconstruct normal samples (Gong et al., 2019; Zhou and Paffenroth, 2017; Zhao et al., 2017; Liu et al., 2021), and generative models like GANs (Schlegl et al., 2017; Sabuhi et al., 2021), normalizing flows (Gudovskiy et al., 2022), and diffusion models (Livernoche et al., 2023). However, several studies (Kirichenko et al., 2020; Zhang et al., 2021) have shown that generative models trained to match the marginal data distribution can fail to identify out-of-distribution or context-specific anomalies, as they may assign high likelihood to atypical but structurally plausible inputs.
To address this, self-supervised learning methods (Golan and El-Yaniv, 2018; Wang et al., 2019b; Chen et al., 2020; Tack et al., 2020) recast anomaly detection as an auxiliary classification task, such as predicting data augmentations or transformations. While effective in image domains, extending such techniques to tabular or categorical data is non-trivial (Bergman and Hoshen, 2020; Qiu et al., 2021; Shenkar and Wolf, 2022), and performance can degrade when the auxiliary task is misaligned with anomaly structure.
Orthogonal to these, subspace anomaly detection methods aim to discover subsets of features or projections where anomalies become more apparent. Classical examples include axis-parallel subspace approaches (Kriegel et al., 2009), projection pursuit, and local subspace analysis (Aggarwal and Yu, 2013; Sathe and Aggarwal, 2016). These methods, including randomized hashing (Sathe and Aggarwal, 2016), avoid modeling the full joint distribution by seeking dimensions that isolate outliers. However, they typically treat all features equally and do not distinguish between context and behavioral attributes.
In contrast, context-aware or conditional anomaly detection explicitly models how behavioral features deviate from expected values given a specific context (Song et al., 2007; Stein et al., 2020). For instance, CADENCE (Amin et al., 2019) models the conditional probability of discrete events within high-cardinality logs, and cross-linked VAEs (Shulman et al., 2020) disentangle context and content in anomaly scoring. Similarly, kernel mixture networks (Kim and Kim, 2023) estimate flexible context-conditioned distributions in sensor data. These approaches highlight the benefits of modeling conditional distributions over global joint distributions in heterogeneous environments, but they often assume that the context is known a priori, or restrict themselves to domains like time-series or event logs.
Our work builds on this literature in three key ways. First, we target tabular data—a setting with less structure and no predefined context—making the task of identifying and modeling contextual dependencies significantly more challenging. Second, unlike methods that rely on dual encoders or complex density estimators, we propose a lightweight yet effective Conditional Wasserstein Autoencoder (CWAE) that models $P(\text{content}|\text{context})$ using a simple reconstruction loss and MMD regularization. Third, we introduce a practical, unsupervised context selection strategy based on early validation loss, allowing our method to automatically discover useful contextual features without supervision. We show that this strategy improves anomaly detection performance across a diverse benchmark suite, demonstrating that meaningful contextual structure often exists—even when not explicitly labeled—and can be effectively leveraged with minimal complexity.
3. Preliminaries
Consider a random variable $\mathbf{X}\in\mathbb{R}^{d}$ representing data with $d$ features. We assume that $\mathbf{X}$ follows distribution $P(\mathbf{X})$. The anomaly detection problem aims to determine whether $\mathbf{X}$ is typical with respect to $P(\mathbf{X})$ or anomalous.
Definition (Context-Conditional Anomaly Detection) Let $\mathbf{X}=(\mathbf{C},\mathbf{Y})\in\mathbb{R}^{k}\times\mathbb{R}^{d-k}$ where $\mathbf{C}$ represents $k$-dimensional context features and $\mathbf{Y}$ represents (dk)(d-k)-dimensional content features. Given a training dataset $\mathcal{D}=\{(\mathbf{c}_{i},\mathbf{y}_{i})\}_{i=1}^{n}$ drawn from the distribution of normal events, our objective is to learn a score function
$S:\mathbb{R}^{k}\times\mathbb{R}^{d-k}\rightarrow\mathbb{R}^{+}$
that assigns an anomaly score $S(\mathbf{X})$ to each data point $\mathbf{X}$. Then based on this anomaly score, we set a threshold $\tau$ so that points with scores $S(\mathbf{X})$ exceeding $\tau$ are labeled as anomalies.
In traditional anomaly detection research, a common approach is to model the joint probability distribution over the entire feature space and to construct a single decision boundary to separate normal and anomalous behaviors. For instance, suppose $\mathbf{X}=\{A,B,C,D\}$ with four features, and we assume $\mathbf{X}$ follows a joint distribution $P(\mathbf{X})=P(A,B,C,D)$. We estimate $P(\mathbf{X})$ from the training data and then derive an anomaly score directly from the learned distribution. Typically, observations with lower probability under $P(\mathbf{X})$ are considered more likely to be anomalous. Formally:
$\text{Score}(\mathbf{X})=S(P(\mathbf{X})),$
or some other monotonic transformation of $P(\mathbf{X})$, i.e., $\log P(\mathbf{X})$. To determine whether a new data point $\mathbf{X}$ is anomalous, the model compares the score against a predetermined threshold $\tau$:
$\text{Label}(\mathbf{X})=\begin{cases}\text{Normal}&\text{if }\text{Score}(\mathbf{X})\leq\tau\\[6.0pt] \text{Anomalous}&\text{if }\text{Score}(\mathbf{X})>\tau.\end{cases}$
Here the single decision boundary $\tau$ is derived from training data, i.e., the maximum anomaly score.
A model using $P(\mathbf{X})$ for anomaly detection without conditional modeling relies on an implicit assumption of homogeneity across different contexts. However, this assumption may fail in practice, particularly when dealing with large-scale and complex datasets, as discussed in Section 1. If the context significantly influences the anomaly patterns, this assumption introduces noise into the learned boundary. For such scenarios, a single decision boundary derived from the joint distribution might not adequately capture the diverse anomaly patterns. Next we introduce a contextual conditioning approach that effectively isolates the variability across different contexts, enabling the model to learn more precise boundaries.
3.1. Context-Conditional Anomaly Detection
To address the heterogeneous nature of the problem, we hypothesize that a subset of features—referred to as optimal context features—captures the greatest diversity of anomaly patterns across their unique values, thereby defining distinct contexts. For simplicity, we initially focus on a single feature that exhibits this maximal diversity, denoted as the context feature, while treating the remaining features as content features to represent anomaly patterns. This choice does not affect the subsequent theoretical results, which naturally extend to multiple context features.
As established in our definition above, we denote $\mathbf{C}$ as $k$-dimensional context features and $\mathbf{Y}$ as (dk)(d-k)-dimensional content features, where $k<d$. Given a training dataset $\mathcal{D}=\{(\mathbf{c}_{i},\mathbf{y}_{i})\}_{i=1}^{n}$ drawn from the distribution of normal events, our objective is to learn a score function
$S:\mathbb{R}^{k}\times\mathbb{R}^{d-k}\rightarrow\mathbb{R}^{+}$
where $S(\mathbf{c},\mathbf{y})$ outputs an anomaly score for the context and content observation $(\mathbf{c},\mathbf{y})$. Unlike the conventional approach that models $P(\mathbf{X})$, we now consider the conditional distribution $P(\mathbf{Y}\mid\mathbf{C})$. By focusing on $P(\mathbf{Y}\mid\mathbf{C}=\mathbf{c})$, we may define context-wise thresholds $\tau_{\mathbf{c}}$ for each $\mathbf{c}\in\text{supp}(\mathbf{C})$.
By leveraging this conditional structure, the model learns a separate decision boundary for each value $\mathbf{c}$ that $\mathbf{C}$ can take. In other words, it adapts the anomaly detection rule to the specific subspace partitioned by each context value. Thus, instead of relying on a single threshold $\tau$, we can define a context-dependent threshold $\tau_{\mathbf{c}}$.
The resulting context-conditioned anomaly detection rule becomes:
$\text{Label}(\mathbf{c},\mathbf{y})=\begin{cases}\text{Normal},&\text{if }S(\mathbf{c},\mathbf{y})\leq\tau_{\mathbf{c}},\\[6.0pt] \text{Anomalous},&\text{if }S(\mathbf{c},\mathbf{y})>\tau_{\mathbf{c}}.\end{cases}$
This approach allows the model to more accurately profile the true boundaries of normal behavior by accounting for contextual variability, rather than relying on a single, global threshold.
4. On the Advantages of Contextual Modeling
Conceptually, our context-conditioned approach parallels discriminative logic in the classic generative–discriminative framework. While generative methods strive to capture the joint distribution of inputs and labels (analogous to content and context in our framework), they can be error-prone if assumptions about the data-generating process are misspecified. By contrast, a conditional or discriminative approach sidesteps modeling the context distribution altogether, focusing solely on how content depends on context. This distinction is particularly salient when context variables (e.g., userIDs) have high cardinality, since modeling $P(\mathbf{C})$ or the full joint distribution $P(\mathbf{Y},\mathbf{C})$ becomes prohibitively complex, risking a curse of dimensionality.
Such complexity motivates Vapnik’s principle (Vapnik, 1995), which asserts that one should not solve a more general, and potentially more difficult, problem than necessary to meet the ultimate predictive goal. By focusing on $P(\mathbf{Y}\mid\mathbf{C})$, our method reduces the risk of model misspecification surrounding context variables and streamlines parameter estimation in high-dimensional spaces. In practice, this can lead to more robust and data efficient performance, which is an advantage also highlighted by extensive generative vs. discriminative comparisons in the classification literature. We illustrate this in more detail in the following subsections.
Furthermore, we illustrate the advantage of conditional modeling from the perspective of variance reduction by the conditioned context variable. We argue that with an optimal choice of context variable, the conditional variance given each context would be smaller, resulting in less spread among data points, making deviations or anomalies more pronounced and easier to detect.
4.1. Model Misspecification Risks
Several theoretical and empirical studies substantiate these advantages for conditional or discriminative approaches. (Ng and Jordan, 2001) provides a rigorous comparison of Logistic Regression versus Naïve Bayes, demonstrating that although the generative model can dominate in very low-data regimes (due to stronger assumptions reducing variance), the discriminative model typically surpasses it once the sample size becomes sufficiently large. This phenomenon occurs precisely because a discriminative model directly optimizes the conditional likelihood pertinent to the task—akin to estimating $P(\mathbf{Y}\mid\mathbf{C})$ rather than the full joint distribution. The asymptotic analysis by (Liang and Jordan, 2008) further clarifies how misspecification in modeling $P(\mathbf{C})$ can degrade a generative model’s performance, while a discriminative estimator that only captures $P(\mathbf{Y}\mid\mathbf{C})$ remains robust. (Lasserre et al., 2006) shows how purely generative assumptions can degrade classification accuracy when they are violated, and how incorporating a discriminative component can improve robustness and prediction quality.
4.2. Curse of Dimensionality
In the high-dimensional context space, (Hastie et al., 2009) discusses how focusing on the conditional distribution helps avoid unnecessary parameter explosion, while (Bishop, 2006) and (Murphy, 2012) provide extensive treatments of both paradigms, showing that discriminative techniques can be more robust to structural mismatches in the data generation process. These results generalize beyond simple supervised learning to broader scenarios such as one-class classification or semi-supervised training, where focusing on the part of the distribution necessary for the objective—be it anomaly detection or prediction—can reduce parameter complexity and improve predictive accuracy. Hence, in contexts where one can define a meaningful ”background” subset of variables, instead of learning the full joint distribution, conditioning on them directly can provide both practical and theoretical advantages in model efficiency. By extension, in high-dimensional scenarios, focusing solely on $P(\mathbf{Y}\mid\mathbf{C})$ avoids the exponential blow-up in complexity that can come with modeling or sampling from $P(\mathbf{C})$. Consequently, conditional modeling can confer both practical and theoretical gains in efficiency and accuracy, making it a compelling strategy whenever a meaningful separation between context and content variables is possible.
4.3. Variance Reduction
In this section, we compare joint vs. conditional anomaly detection from a variance‐reduction standpoint. Let
$\mathbf{Y}\in\mathbb{R}^{d-k}\quad\text{(content features)},\qquad\mathbf{C}\in\mathbb{R}^{k}\quad\text{(context features)}.$
A joint model learns the full density $P(\mathbf{Y},\mathbf{C})$, however, the marginal
$P(\mathbf{Y})=\int P(\mathbf{Y},\mathbf{C})\,\mathrm{d}\mathbf{C}=\int P(\mathbf{Y}\mid\mathbf{C})\,P(\mathbf{C})\,\mathrm{d}\mathbf{C}$
mixes together data from all contexts.
By contrast, a conditional approach directly learns $P(\mathbf{Y}\mid\mathbf{C})$ and scores each sample within its own context slice. The key theoretical insight is the law of total variance:
$\operatorname{Var}(\mathbf{Y})$ $=\underbrace{\mathbb{E}_{\mathbf{C}}\bigl{[}\operatorname{Var}(\mathbf{Y}\mid\mathbf{C})\bigr{]}}_{\begin{subarray}{c}\text{average "within-context"}\\ \text{variance}\end{subarray}}+\,\,\,\,\,\underbrace{\operatorname{Var}_{\mathbf{C}}\!\bigl{(}\mathbb{E}[\mathbf{Y}\mid\mathbf{C}]\bigr{)}}_{\begin{subarray}{c}\text{"between-context"}\\ \text{variance}\end{subarray}}.$
A joint detector must accommodate both terms, whereas a context‐conditional detector only needs to cope with $\operatorname{Var}(\mathbf{Y}\mid\mathbf{C}=\mathbf{c})$, which is typically much smaller than the total $\operatorname{Var}(\mathbf{Y})$.
So ideally if we can find contexts to partition the data so that
$\operatorname{Var}(\mathbf{Y}\mid\mathbf{C}=\mathbf{c})\ll\operatorname{Var}(\mathbf{Y})$
holds for most $\mathbf{c}\in\text{supp}(\mathbf{C})$, conditioning makes our anomaly detectors both sharper and more sensitive to small deviations from normal, context‐specific patterns. For reconstruction based models, this variance reduction yields the following practical benefits:
  1. (1)
    Higher signal‐to‐noise ratio. Normal samples concentrate more tightly around their reconstructions, making anomalous deviations stand out. We can choose a threshold $\tau_{\mathbf{c}}$ for each context $\mathbf{c}$ based on the smaller $\operatorname{Var}(\mathbf{Y}\mid\mathbf{C}=\mathbf{c})$, instead of a single conservative $\tau$ for the full marginal distribution.
  2. (2)
    More stable and efficient learning. Each conditional model only needs to capture a homogeneous slice of the data, leading to more stable gradient updates, faster convergence, and fewer mixed‐mode reconstruction errors, as the model doesn’t need to accommodate conflicting patterns from different contexts simultaneously.
5. Modeling Conditional Distribution with CWAE
In this work, we introduce the Contextual Wasserstein Auto-Encoder (CWAE), a straightforward autoencoder-based method for modeling context-conditional distributions in anomaly detection. CWAE is designed as a proof-of-concept to showcase how incorporating contextual information can improve detection performance. Although not intended as an optimized or state-of-the-art solution, it offers a clear and interpretable framework for illustrating the advantages of context-aware learning in anomaly detection.
5.1. CWAE Model Structure
The CWAE model is designed to learn the conditional distribution of input features given contextual features. This section outlines the objective, architecture, and loss function.
5.1.1. Model Objective
CWAE aims to conditionally reconstruct input content features based on associated context features. By learning a latent representation that combines content and context, the model identifies anomalies through reconstruction loss, where higher reconstruction loss indicates a greater likelihood of anomalous behavior. This framework allows CWAE to adapt to varying contextual conditions, which is critical for datasets with complex or diverse anomaly patterns.
5.1.2. Model Architecture
Figure 1. Architecture of the CWAE model utilized in the experiments.
CWAE consists of three main components: embedding layers, an encoder, and a decoder.
Embedding Layers. Two sets of embedding layers are employed—one for context features and another for content features. Each embedding layer is randomly initialized, with the first dimension corresponding to the number of unique feature values plus one (to handle unseen values) and the second dimension representing the embedding vector size.
Encoder. The encoder takes as input the concatenated embeddings of context and content features. This input is processed through fully connected (FC) layers to extract a latent representation. To retain contextual information, the output of these layers is concatenated with the original context embeddings to form the final latent representation.
Decoder. The decoder reconstructs the content features from the latent representation. A series of FC layers are used to map the latent representation back to the original feature space.
5.1.3. Loss Function
The loss function combines reconstruction accuracy with regularization to ensure effective learning. Formally:
$\mathcal{L}_{\text{cwae}}$ $=\sum_{i=1}^{d}\text{CE}(\mathbf{Y}_{i},\widehat{\mathbf{Y}}_{i})$
(1) $\quad+\lambda\cdot\text{MMD}\Big{(}P(z),Q(z\mid X=\mathbf{Y},\mathbf{C})\Big{)},$
where $\mathbf{Y}$ and $\mathbf{C}$ represent context and content features, respectively; $\text{CE}(\mathbf{Y}_{i},\widehat{\mathbf{Y}}_{i})$ represents the cross-entropy reconstruction loss for content features and d is the dimension of content $\mathbf{Y}$; $\lambda$ is a weighting hyperparameter for the MMD term; $\text{MMD}\left(P(z),Q(z\mid X)\right)$ quantifies the maximum mean discrepancy between the latent distribution and a predefined prior (e.g., Gaussian). The reconstruction loss ensures that input features are accurately reconstructed, while the MMD regularization term aligns the latent distribution with the desired prior. The combined loss facilitates effective training and ensures the model captures meaningful contextual patterns.
5.2. Training and Inference
Training. During training, tabular features are first mapped to indices, resulting in integer vectors as input. These indices are passed through embedding layers, which are updated during training. The encoder and decoder work together to reconstruct the content features, and the total loss is optimized to minimize reconstruction error and regularization discrepancy.
Inference. During inference, anomaly detection is performed by calculating the reconstruction loss for each instance. Samples with reconstruction loss exceeding the context-specific threshold (as defined in Section 6.1.3) are classified as anomalies. This approach allows the model to adapt to varying contextual conditions while maintaining robust detection performance.
5.3. Selecting the Right Feature for Conditioning
Selecting an appropriate context feature is one of the most critical and challenging aspects of contextual anomaly detection. The chosen context fundamentally shapes how the model interprets patterns in the data and can have a substantial impact on detection performance. As shown in Tables 3, 4, and 5, contextual models consistently outperform their non-contextual counterparts when the context is well chosen, underscoring its central role in effective anomaly detection. Achieving strong results with CWAE therefore requires a principled and effective strategy for selecting a context feature that is likely to yield good performance.
We propose to frame context selection as a bilevel optimization problem (Sinha et al., 2018), where the goal is to optimize for the joint data likelihood $P(\mathbf{Y},\mathbf{C})=P(\mathbf{Y}\mid\mathbf{C})\,P(\mathbf{C})$. In this setting, both the model parameters and the context feature are treated as variables to optimize. The inner loop learns the model weights for a given context, while the outer loop searches for the context feature that yields the best generalization performance.
Naively training each conditional model to convergence is accurate, but can be prohibitively expensive. In practice, we only train the model for one epoch, and evaluate the model with validation loss afterwards. We observe that this approach serves as an effective proxy for generalization quality. Numerous prior works have demonstrated the effectiveness of leveraging an early-trained model for other tasks. Similar ideas have been applied for neural architecture search (Liu et al., 2019; Bender et al., 2018), neural network model pruning (You et al., 2020), dataset pruning (Paul et al., 2021) and learning with noisy labels (Liu et al., 2020; Zhang and Sabuncu, 2018). Prior works have also justified both empirically and theoretically the importance of the first few epochs for generalization of neural networks (Achille et al., 2017; Zhang et al., 2016; Frankle et al., 2020).
To ensure consistency across models with different probabilistic factorizations—e.g., $P(\mathbf{Y}\mid\mathbf{C})$ vs. $P(\mathbf{Y},\mathbf{C})$—we evaluate all models using their joint log-likelihood. For conditional models, we apply the identity $P(\mathbf{Y},\mathbf{C})=P(\mathbf{Y}\mid\mathbf{C})\,P(\mathbf{C})$, where $P(\mathbf{C})$ is estimated empirically from the training data. This formulation also penalizes rare context values via the $\log P(\mathbf{C})$ term, favoring parsimonious and meaningful conditioning. For simplicity, we only condition on one feature at a time as the context feature. As such, $P(\mathbf{C})$ can be easily obtained from the training data.
Putting everything together we present the proposed procedure for selecting context features in Algorithm 1. This data-driven selection method avoids heuristic filtering and offers a principled framework for identifying useful context features.
Algorithm 1 Context Selection via Validation Loss Minimization
Given candidate context features $\mathcal{K}$, training set $\mathcal{D}_{\text{train}}$, validation set $\mathcal{D}_{\text{val}}$, and model $\mathcal{M}$:
  1.  1.
    Outer Loop (Context Optimization):
    • For each $\mathbf{C}\in\mathcal{K}$:
      1.  a.
        Initialize model $\mathcal{M}$ conditioned on context $\mathbf{C}$
      2.  b.
        Inner Loop (Model Training): Train $\mathcal{M}$ on $\mathcal{D}_{\text{train}}$ to optimize weights $w$ for 1 epoch
      3.  c.
        Compute joint validation loss $\mathcal{L}_{\text{val}}=-\log P(\mathbf{Y},\mathbf{C})$ on $\mathcal{D}_{\text{val}}$
  2.  2.
    Select context $\mathbf{C}^{*}$ with minimal validation loss
Return final context $\mathbf{C}^{*}$ and trained model $\mathcal{M}(\mathbf{Y}\mid\mathbf{C}^{*})$
Figure 2 illustrates this process on the Census dataset using CWAE. We show the validation loss curves for the top and bottom three candidate context features, based on performance after a single epoch. The feature detailed_occupation_recode results in the lowest loss and is therefore selected.
Using this selection method, we achieve performance equal to or better than the non-contextual baseline in all but two datasets. Figure 3 shows the improvement in AUCROC for CWAE when using the selected context feature, compared to its non-contextual counterpart. In cases where the no-context option yields the best validation loss, it is retained as the final choice (marked in orange). Final context selections are summarized in Table 3 under the Context column.
Figure 2. Validation loss curves for CWAE on the Census dataset, conditioned on different context features.
Figure 3. AUCROC improvement achieved by applying Algorithm 1 to select context for CWAE across datasets.
5.4. Advantages of the Wasserstein Approach
As outlined in Section 4, effective conditional modeling of $P(\mathbf{Y}\mid\mathbf{C})$ benefits from architectures that are both stable and efficient at inference. While Variational Autoencoders (VAEs) (Sohn et al., 2015) have been widely used for generative modeling, their non-deterministic encoders require multiple latent samples at inference to obtain stable anomaly scores, increasing computational cost and introducing score variability.
Wasserstein Autoencoders (WAEs) replace the KL-divergence regularization in VAEs with an optimal transport–based term such as Maximum Mean Discrepancy (MMD), enabling fully deterministic encoders without the overfitting issues of unregularized autoencoders. Deterministic encoding removes the need for Monte Carlo sampling, making scoring faster and more stable—aligning with the robustness and efficiency goals discussed in Section 4.
In this work, we instantiate CWAE with a deterministic encoder–decoder pair. This design choice reduces inference latency, improves score consistency, and maintains a well-regularized latent space, providing practical benefits over the Conditional VAE (CVAE) baseline while preserving the ability to model complex conditional distributions.
6. Experiments
6.1. Experiment Setup
6.1.1. Dataset Selection
To evaluate contextual learning for anomaly detection, we selected datasets spanning diverse domains, including finance, cybersecurity, demographics, and network intrusion detection. Each dataset is designed to facilitate the study of anomaly detection with contextual learning, where anomalies are defined by specific criteria such as customer behavior, malicious events, or demographic thresholds.
We started with datasets introduced in the AD literature, such as those discussed in (Han et al., 2022), and expanded our selection to include additional datasets originally used for multi-class classification. Collectively, we have compiled a suite of eight datasets for comprehensive empirical evaluation, which includes the Bank Marketing dataset (bank) (Moro et al., 2012), the Beth Cybersecurity dataset (beth) (Highnam et al., 2021), the Census dataset (census) (noa, 2000b), the CMC dataset (cmc) (Tjen-Sien Lim, 1999), the KDD dataset (kdd) (noa, 1999a), the LANL dataset (lanl) (Turcotte et al., 2019), the Solar Flares dataset (sf) (Unknown, 1989) and the Spotify dataset (spotify) (Choksi, 2021). Details about the datasets can be found in Table 1. These datasets vary significantly in size, feature count, and anomaly prevalence, providing a comprehensive foundation for robust evaluation.
6.1.2. Data Complexity Ranking
Understanding dataset complexity is essential for ensuring that contextual learning methods are evaluated across a broad spectrum of dataset difficulties. By considering complexity, we can better assess how contextual learning behaves in both simpler and more challenging anomaly detection scenarios. This perspective allows us to examine whether the benefits of contextual learning hold consistently across varying levels of difficulty. To capture these complexities, we adopt several metrics from AD literature (Pang et al., 2021a), tailored to reflect the specific characteristics of our datasets. These metrics are defined as:
  • Value Coupling Complexity ($K_{\text{vcc}}$): Quantifies the impact of similarity among outliers (homophily) on detection effectiveness. Higher $K_{\text{vcc}}$ values indicate greater complexity, as coupled outliers introduce noise and correlations that hinder detection.
  • Heterogeneity of Categorical Distribution ($K_{\text{het}}$): Measures the diversity in the frequency distribution of dominant categories across features. A higher $K_{\text{het}}$ value signifies increased difficulty in identifying outliers due to the varied distribution of categorical modes.
  • Outlier Inseparability ($K_{\text{ins}}$): Represents the difficulty of separating outliers from normal data points. It ranks features based on the inverse frequency of their values, with higher $K_{\text{ins}}$ values suggesting that anomalies are less distinguishable from normal objects.
  • Feature Noise Level ($K_{\text{fnl}}$): Captures the proportion of features where noise causes outliers to exhibit frequencies similar to or higher than normal data points. Higher $K_{\text{fnl}}$ values indicate increased detection difficulty due to noise obscuring distinctions.
Table 2 provides a comprehensive breakdown of dataset complexity across the four distinct metrics: $K_{\text{vcc}}$, $K_{\text{het}}$, $K_{\text{ins}}$, and $K_{\text{fnl}}$. For each dataset, we report the raw metric value, its min-max scaled version, and the corresponding rank. The overall complexity score (Avg Scaled) is computed as the average of the scaled scores across all four metrics. This aggregate score serves as a unified measure of dataset difficulty. The final ranking reflects this overall complexity, where higher scores correspond to more challenging datasets for anomaly detection, and lower rank values indicate relatively easier datasets.
Dataset Name Number of Features Size Number of Anomalies Anomaly Ratio Avg Cardinality bank 11 41,188 4,640 11.27% 5 beth 11 1,141,078 158,432 13.88% 35,154 census 38 299,285 18,568 6.20% 17 cmc 8 1,473 29 1.97% 3.13 kdd 7 1,014,535 440 4.51% 12 lanl 16 2,542,727 5,971 0.23% 3,191 sf 11 1,066 43 4.03% 3.73 spotify 17 113,550 2,412 2.12% 4,615
Table 1. Overview of datasets with descriptive statistics.
$K_{vcc}$ $K_{het}$ $K_{ins}$ $K_{fnl}$ Overall Complexity
Dataset Score Scaled Rank Score Scaled Rank Score Scaled Rank Score Scaled Rank Avg Scaled Rank
bank 0.210 0.196 5 2.015 0.016 4 0.343 0.775 3 1.000 1.000 1 0.497 2
beth 0.916 1.000 1 2.929 0.037 3 0.016 0.032 7 0.179 0.179 5 0.312 4
census 0.450 0.469 3 3.500 0.049 2 0.238 0.536 4 0.286 0.286 4 0.335 3
cmc 0.038 0.000 8 1.579 0.007 6 0.348 0.786 2 0.000 0.000 8 0.198 7
kdd 0.055 0.019 7 1.278 0.000 8 0.159 0.357 6 0.500 0.500 2 0.219 6
lanl 0.369 0.377 4 1.939 0.015 5 0.002 0.000 8 0.009 0.009 7 0.100 8
sf 0.124 0.098 6 1.564 0.006 7 0.176 0.395 5 0.500 0.500 2 0.250 5
spotify 0.500 0.526 2 46.264 1.000 1 0.442 1.000 1 0.123 0.123 6 0.662 1
Table 2. Dataset complexity metrics with raw scores (Score), scaled scores (Scaled), and ranks (Rank) reported for each. Higher scores indicate greater complexity, while lower ranks correspond to higher complexity.
Comparing Tables 1 and 2 shows that our complexity metrics capture challenges that basic statistics miss. For instance, spotify emerges as the most complex dataset despite appearing ordinary in size, feature count, and cardinality, highlighting the role of structural and contextual patterns in driving difficulty.
Conversely, datasets with large size or high cardinality, such as beth and lanl, rank lower in complexity, suggesting that these properties alone do not necessarily make anomaly detection harder. Overall, the complexity metrics offer a more nuanced basis for evaluating contextual learning across datasets of varying difficulty.
6.1.3. Contextual Thresholding for AUCROC Calculation
Traditional AUCROC calculation typically involves grid searching over a single, global decision threshold. In the no-context setting, this is straightforward since all samples share the same threshold. However, in contextual learning, using a single global threshold can be suboptimal because score distributions may vary significantly across context groups. To address this, we adopt a per-context thresholding approach while ensuring that resulting AUCROC values remain directly comparable to the baseline.
Given a training set with a particular context feature selected, we define the maximum training loss within each context group as its contextual threshold $H_{\mathbf{C}_{i}}$. For the test set, each sample’s anomaly score is normalized by its corresponding contextual threshold, producing a contextual ratio $R_{\mathbf{C}_{i}}=\frac{\text{anomaly score}}{H_{\mathbf{C}_{i}}}$. This normalization rescales scores so that they are on the same relative scale across different contexts, effectively reducing the per-group thresholding problem to a single normalized thresholding problem.
We then compute the AUCROC by grid searching over 100 evenly spaced thresholds across the range of $R_{\mathbf{C}_{i}}$. For example, step $k$ corresponds to a threshold of $0.01\cdot k\cdot\max(R_{\mathbf{C}_{i}})$. Since both the contextual and no-context settings ultimately apply thresholding to a single unified score distribution (global anomaly scores in the baseline, normalized contextual ratios in the contextual case), the resulting AUCROC values are directly comparable across all model configurations.
6.1.4. Benchmark Models
To effectively benchmark our proposed model, we selected six SOTA anomaly detection models from the DeepOD package (Xu, 2024): DSVDD (Ruff et al., 2018), RDP (Wang et al., 2019a), RCA (Liu et al., 2021), ICL (Shenkar and Wolf, 2022), DIF (Xu et al., 2023a), and SLAD (Xu et al., 2023b). Based on our research, these models represent leading techniques in the anomaly detection literature. The DeepOD package provides comprehensive implementations of these methods within a unified testbed, enabling consistent evaluation. Each model was tested using the default hyperparameter settings provided by DeepOD, ensuring comparability across experiments. All models were implemented without incorporating contextual considerations for anomaly detection.
In addition to the DeepOD-based models, we include the DTE (Livernoche et al., 2023) model, implemented from the available paper details. As one of the most recent diffusion-based approaches to anomaly detection, DTE introduces a fundamentally different architecture that broadens the diversity of our baseline set. We also evaluate the WAE (Tolstikhin et al., 2019) model as a non-contextual baseline. WAE shares the same architecture as our proposed CWAE but treats all features uniformly as content, without distinguishing contextual variables, making it a strong reference point for isolating the effects of contextual modeling.
6.2. Performance Evaluation
Dataset Scaled Complexity DSVDD RDP RCA ICL DIF SLAD DTE WAE CWAE Context
bank 0.497 0.455 0.582 0.649 0.519 0.580 0.464 0.582 0.654 0.687 loan
beth 0.312 0.895 0.998 0.997 0.953 0.988 0.995 0.999 0.996 0.996 no_ctx
census 0.335 0.443 0.629 0.701 0.637 0.669 0.673 0.410 0.676 0.743 detailed_occupation_recode
cmc 0.198 0.610 0.566 0.760 0.530 0.675 0.713 0.699 0.702 0.751 Husbands_education
kdd 0.219 0.615 0.924 0.880 0.501 0.928 0.939 0.883 0.674 0.639 is_guest_login
lanl 0.100 1.000 0.948 0.867 1.000 0.914 0.978 1.000 0.999 0.999 SubjectLogonID
sf 0.250 0.579 0.811 0.800 0.731 0.827 0.267 0.731 0.678 0.895 C-class_flares_production_by_this_region
spotify 0.662 0.403 0.537 0.501 0.585 0.480 0.531 0.431 0.697 0.665 loudness
AVG 0.625 0.749 0.769 0.682 0.758 0.695 0.717 0.760 0.797
Table 3. AUCROC scores, scaled complexity scores, and the selected context feature for all models across datasets. Bold and underlined values indicate the best AUCROC scores.
Dataset DSVDD RDP RCA ICL DIF SLAD DTE WAE CWAE
bank 9 4 3 7 6 8 4 2 1
beth 9 2 3 8 7 6 1 4 4
census 8 7 2 6 5 4 9 3 1
cmc 7 8 1 9 6 3 5 4 2
kdd 8 3 5 9 2 1 4 6 7
lanl 1 7 9 1 8 6 1 4 4
sf 8 3 4 5 2 9 5 7 1
spotify 9 4 6 3 7 5 8 1 2
AVG 7.375 4.75 4.125 6 5.375 5.25 4.625 3.875 2.75
Table 4. Model ranks for all datasets. Bold and underlined values indicate the best (lowest) rank.
In this section, we present the CWAE results obtained using the selection method described in Section 5.3. Tables 3 and 4 report the corresponding AUCROC scores and their rankings, respectively.
(1) CWAE achieves top overall performance. We evaluate model performance using two metrics: AUCROC and average rank. Ranks are computed per dataset, then averaged across all datasets to offer a balanced view that reduces the influence of outliers. CWAE attains the best overall performance, with the highest average AUCROC (0.797) and the lowest average rank (2.75) among all models.
(2) Contextual learning improves over non-contextual baselines. CWAE consistently outperforms its non-contextual counterpart WAE, which shares the same architecture but omits context features. While WAE achieves an average AUCROC of 0.760 (rank 3.875), CWAE improves upon it across both metrics, highlighting the benefit of incorporating context in anomaly detection. Compared to other leading non-contextual baselines of different architectures—RCA (0.769, 4.125), DIF (0.758, 5.375), RDP (0.749, 4.75), and DTE (0.717, 4.625)—CWAE remains the strongest overall.
(3) CWAE excels across datasets of varying complexity. A dataset-wise breakdown reveals CWAE’s flexibility and robustness:
  • spotify (high complexity): WAE achieves the best AUCROC (0.697), but CWAE is competitive at 0.665 and outperforms several baselines, including ICL (0.585) and RDP (0.537). This illustrates contextual modeling’s effectiveness in challenging data regimes.
  • bank and census (moderate complexity): CWAE attains the top AUCROC on both datasets (0.687 and 0.743), surpassing RCA and WAE. This shows strong generalization to structured real-world tabular data.
  • lanl (low complexity): CWAE attains a near-perfect AUCROC of 0.999, just behind DSVDD, ICL, and DTE at 1.0, showing that generalization is not limited to high-complexity regimes.
  • beth (high cardinality): On this large and high-cardinality dataset, CWAE (0.996) matches WAE and trails only slightly behind DTE (0.999), demonstrating scalability.
(4) Consistency distinguishes CWAE from other models. CWAE ranks first or second on five of the eight datasets, spanning a spectrum of complexity levels. In contrast, models like RCA, SLAD, or DTE achieve top results only on one or two dataset and suffer from inconsistency elsewhere, as reflected by their higher average ranks.
These results affirm the core message of this work: context matters. When paired with effective context selection, CWAE demonstrates scalable, generalizable, and state-of-the-art anomaly detection performance. The observed gains are consistent across a variety of domains and not limited to any single dataset characteristic, underscoring the broader value of integrating contextual information.
7. Ablation Studies
Dataset Best SOTA WAE CWAE CWAE Best CWAE Context CWAE Best Context
AUCROC Rank AUCROC Rank AUCROC Rank AUCROC Rank
bank 0.649 4 0.654 3 0.687 2 0.695 1 loan education
beth 0.999 1 0.996 3 0.996 3 0.997 2 no_ctx argsNum
census 0.701 3 0.676 4 0.743 1 0.743 1 detailed_occupation_recode detailed_occupation_recode
cmc 0.760 2 0.702 4 0.751 3 0.787 1 Husbands_education Contraceptive_method_used
kdd 0.939 1 0.674 3 0.674 3 0.872 2 no_ctx service
lanl 1.000 1 0.999 3 0.999 3 1.000 1 SubjectLogonID AuthenticationPackage
sf 0.827 3 0.678 4 0.895 2 0.906 1 C-class_flares_production_by_this_region X-class_flares_production_by_this_region
spotify 0.585 4 0.697 2 0.665 3 0.711 1 loudness valence
AVG 0.808 2.375 0.760 3.25 0.801 2.5 0.839 1.25
Table 5. Comparison of AUCROC scores and ranks for WAE, CWAE, CWAE Best, and the best SOTA model per dataset. Bold and underlined values indicate the best AUCROC and best (lowest) rank.
7.1. Benefits of Contextual Thresholding
Contextual anomaly detection models allow for context-specific thresholds, which offer a more nuanced decision boundary than a global threshold approach. This is particularly relevant in domains where the distribution of normal behavior varies significantly across groups defined by context features.
Figure 4 visualizes this idea by plotting the thresholds learned by CWAE on the bank dataset, conditioned on different context features. Each point corresponds to the threshold for a given group within a context feature. These thresholds $H_{\mathbf{C}_{i}}$ are obtained in the same manner described in Section 6.1.3.
The primary takeaway from this figure is that thresholds differ significantly across groups, confirming our earlier claim that learning thresholds for each conditional distribution is desirable. This level of customization enables the model to detect context-dependent anomalies that may be missed under a single global threshold—particularly if such anomalies lie below the global cutoff but are relatively anomalous within their specific context.
Limitations. It is important to note that comparing raw threshold values across contexts or between contextual and non-contextual models is not strictly valid. The threshold for WAE corresponds to $P(\mathbf{Y})$, while the thresholds for CWAE correspond to $P(\mathbf{Y}\mid\mathbf{C})$. Since the marginal $P(\mathbf{C})$ is not uniform across all datasets or features, and since scaling thresholds by $P(\mathbf{C})$ would alter the decision boundary, no fair normalization is possible without further assumptions.
Thus, the purpose of this visualization is not to support a direct comparison against WAE, but rather to illustrate that contextual models naturally learn different thresholds for different conditions. This leads to a more flexible and adaptive detection policy, consistent with the goals of contextual anomaly detection.
Figure 4. Thresholds for CWAE on the bank dataset, conditioned on different context features. Each point represents a threshold learned under a specific context group.
7.2. Upper Bound Performance with Optimal Context
In addition to evaluating CWAE under algorithmically selected context features, we assess its performance in the optimal context setting—where the context feature yielding the highest AUCROC score for each dataset is selected. This analysis reveals the potential upper bound of CWAE when context is ideally chosen. Table 5 presents the comparison where CWAE Best indicates the optimal context choices.
(1) Optimal context significantly boosts CWAE performance. On average, CWAE Best achieves an AUCROC of 0.839, outperforming the standard CWAE score of 0.797. It ranks first on 6 of the 8 datasets, compared to only 3 for the standard CWAE. This gap underscores the potential benefit of improved or learned context selection methods.
(2) CWAE Best consistently outperforms the non-contextual baseline. Across all datasets, CWAE Best yields higher AUCROC scores than the non-contextual WAE baseline, often by substantial margins. For instance, on the cmc dataset, CWAE Best achieves 0.787 compared to WAE’s 0.702, and on kdd, it reaches 0.872 versus WAE’s 0.674. These improvements reinforce the core hypothesis of this work: that modeling context is a critical component for effective anomaly detection.
(3) CWAE Best surpasses all SOTA models on most datasets. We aggregate the best scores from all individual baselines into a Best SOTA column to establish a performance reference. CWAE Best outperforms the Best SOTA model on 6 out of 8 datasets, including notable gains such as +0.046 on bank, +0.027 on cmc, and +0.079 on sf. On average, CWAE Best ranks first with a rank of 1.25, compared to Best SOTA’s 2.375. Importantly, CWAE achieves this level of performance using a single unified architecture across all datasets, underscoring its value as a flexible, general-purpose anomaly detection model.
(4) CWAE Best performs well across the full range of dataset complexities. Whether the dataset is low-complexity (e.g., lanl) or high-complexity (e.g., spotify), CWAE Best consistently delivers strong results. This suggests that with appropriate context, CWAE is broadly applicable across varied data regimes.
These results further support the importance of context selection. When given the optimal context, CWAE matches or surpasses state-of-the-art performance, demonstrating the potential of context-aware approaches. This suggests that future research should focus on improving context selection methods, including the development of learnable strategies that can consistently achieve performance close to this upper bound in practice.
7.3. Quantifying the Value of Context
A central question in this research is whether incorporating contextual information improves anomaly detection performance not only across datasets of varying complexity, but also different model architectures. To investigate, we ran additional experiments with other conditional models from the AD literature. The selected models are listed below:
  • Conditional Wasserstein Autoencoder (CWAE), described in Section 5.1.
  • Conditional Variational Autoencoder (CVAE) (Pol et al., 2020), which models the conditional distribution $P_{\theta}(\mathbf{Y}\mid\mathbf{C})$ of observations given contextual variables using a deep generative model with learned feature-wise reconstruction variance.
  • Conditional Mixture Density Network (CMDN) (Dai et al., 2025), which learns a neural network–parameterized Gaussian mixture model to capture $P(\mathbf{Y}\mid\mathbf{C})$, enabling mixture parameters to adapt dynamically based on context.
  • Contextual Anomaly Detection using Isolation Forest (CADI) (Yepmo et al., 2024), which extends the isolation forest algorithm with density-aware splits to jointly detect anomalies and explain them relative to local contextual clusters.
All models are implemented with only minor modifications from their original formulations, using the same embedding layers outlined in Section 5.1. For each model–dataset pair, we systematically iterated over all candidate features as potential context, designating exactly one feature as context in each run and treating the remainder as content. Performance was compared to a no-context baseline in which all features are treated as content. This controlled setting enables a direct, interpretable measure of how both context and model architecture contribute to performance improvements.
To quantify the benefits of contextual learning, we report the average improvement in AUCROC across datasets for each model, comparing the best-performing context configuration to its non-contextual counterpart. All contextual models outperform their non-contextual baselines on average, often by a substantial margin. The CWAE model achieves the highest average improvement of +11.69%, followed by CMDN (+7.47%), CVAE (+6.21%), and CADI (+4.18%). These consistent gains support the hypothesis that leveraging context enhances a model’s ability to detect anomalies across architectures, especially when the contextual signal is informative.
Given its strong empirical performance, we selected CWAE as the primary model for further analysis and methodological development. We treat CWAE as a representative upper bound for context-aware models in our framework, using it as a proxy to explore context selection strategies and to benchmark the overall value of contextual information in anomaly detection.
8. Conclusion
In this work, we introduced a context-conditional anomaly detection framework that addresses a fundamental limitation of traditional anomaly detection methods: their reliance on global decision boundaries that fail to capture contextual variations in heterogeneous data. We demonstrated that modeling conditional distributions $P(\mathbf{Y}\mid\mathbf{C})$ rather than joint distributions $P(\mathbf{Y},\mathbf{C})$ offers significant advantages. Our theoretical analysis, grounded in variance decomposition and principles from discriminative learning, showed that conditional modeling effectively isolates intra-context variability while reducing noise from inter-context variations, leading to more precise anomaly detection boundaries. The proposed CWAE model consistently outperformed both its non-contextual baseline (WAE) and SOTA methods on diverse benchmark datasets.
The implications of this work extend beyond the specific CWAE architecture to establish contextual learning as a valuable paradigm for anomaly detection in tabular data. Our findings demonstrate that many real-world datasets contain latent contextual structure that, when properly leveraged, can significantly enhance detection performance compared to traditional global approaches. The framework’s effectiveness across datasets with varying complexity levels highlights its broad applicability. Future research directions include developing more sophisticated context selection algorithms, exploring multi-context conditioning, and extending the framework to other data modalities.
9. Ethical Considerations
This research addresses several important ethical considerations related to anomaly detection systems and their deployment in real-world applications. Our work focuses on improving the accuracy and fairness of anomaly detection through contextual modeling, which has both positive implications and potential risks that must be carefully considered.
Privacy and Data Protection. The datasets used in this study include sensitive information such as demographic data (census), financial records (bank), and cybersecurity logs (beth, lanl). While we used publicly available datasets that have been appropriately anonymized, the contextual approach we propose could potentially be applied to more sensitive data in practice. Organizations implementing contextual anomaly detection must ensure compliance with privacy regulations such as GDPR and implement appropriate data protection measures, including differential privacy techniques where applicable.
Fairness and Bias Mitigation. A key ethical advantage of our contextual approach is its potential to reduce algorithmic bias by explicitly modeling different contexts rather than applying a single global decision boundary. Traditional anomaly detection systems may unfairly flag certain groups as anomalous due to differences in baseline behavior patterns. By conditioning on context features, our approach can learn appropriate norms for different user groups, potentially reducing false positives for underrepresented populations. However, careful selection of context features is crucial to avoid perpetuating existing biases or creating new forms of discrimination. We recommend thorough bias auditing when deploying contextual anomaly detection systems, particularly in high-stakes applications such as fraud detection or security screening.
Transparency and Explainability. While our contextual approach provides more nuanced anomaly detection, it also introduces additional complexity that may impact system interpretability. The use of context-specific thresholds and conditional modeling may make it more difficult for end users to understand why certain decisions were made. Organizations deploying such systems should invest in explainability tools and provide clear documentation of how contextual factors influence anomaly detection decisions, particularly in applications where individuals may be adversely affected by false positive classifications.